OpenStudy (anonymous):

PLEASE HELP :) Find the GCF of 8 and 32.

OpenStudy (tkhunny):

Please find the Prime Factorization of each. The answer is in the Prime Factorization.

OpenStudy (anonymous):

8,32 Factor out the GCF of 8 from each term in the polynomial. (think what are the factors of 8 and 32) 8*(1)=8 8*(4)=32 8(1)+8(4) = 8+32 Factor out the GCF of 8 from 8+32. 8(1+4) The GCF (Greatest Common Factor) is the term in front of the factored expression. 8 <--- GCF

OpenStudy (tkhunny):

Don't think "what are the factors of 8". This may help with THIS problem, but it will not help you in other problems. Think about the Prime Factorization. This will help you on ALL problems - without fail.